A string 50.0 cm long has a mass per unit length equal to 20.0x10^-5 kg/m. To what tension should this string be stretched if it's fundamental frequency is to be 20.0 Hz.


1
Expert's answer
2021-10-18T13:54:24-0400

"l = 50 \;cm = 0.5 \;m \\
m = \frac{M}{l} = 20 \times 10^{-5} \; kg/m \\
l = 20 \;Hz \\
f = \frac{1}{2l} \sqrt{\frac{T}{m}} \\
f^2 = \frac{1}{4l^2} \times \frac{T}{m} \\
T = f^2 \times 4l^2 \times m \\
= 20^2 \times 4 \times 0.5^2 \times 20 \times 10^{-5} \\
= 800 \times 10^{-4} \\
= 8 \times 10^{-2} \;N"